OpenStudy (anonymous):

Please help solve this problem Write an equation for the linear function f in the form f(x) = mx+b that has the given value. f(-4)=4, f(1)=-6

OpenStudy (anonymous):

-4m+ b = 4 m + b = -6 solve these equations, m= -2, b = -4 so, f(x ) = -2x -4

OpenStudy (anonymous):

m is the slope, which can be found by \[(y _{2}-y _{1})/(x _{2}-x _{1})\] Then you can plug m and one of the points in to the equation y=mx+b to solve for b.